OpenStudy (anonymous):

What is the acceleration of a cabinet of mass 45 kilograms if Jake and Ted push it by applying horizontal force of 25 newtons and 18 newtons respectively in the same direction? (Assume a flat, frictionless surface.)

OpenStudy (anonymous):

here total force will be 25 plus 18 newton because both the force are acting in the same direction. and then use the equation force= mass x acceleration or acceleration = force/mass

OpenStudy (anonymous):

can you help me out i dont understand and i really need this question right to pass

OpenStudy (anonymous):

never mind :)

OpenStudy (anonymous):

Total force = 18 + 25 =42 newton, mass =45 kg , therefore acceleration = 42/45